OpenStudy (anonymous):

Graph the function f(x) = (x + 3)3 by hand and describe the end behavior.

OpenStudy (anonymous):

It may helps you http://www.wolframalpha.com/input/?i=f%28x%29+%3D+%28x+%2B+3%293

OpenStudy (mathmale):

@Yakima_Love : I'd suggest you re-write the given function as \[f(x)=\frac{ 1 }{ 3 }x+1.\] This new form is exactly equivalent. To me, this new form is easier to deal with, since it's obvious from inspection that the slope of the line is 1/3 and the y-intercept is 1.

OpenStudy (anonymous):

3 is power? u mean (x+3)^3 ?

OpenStudy (lasttccasey):

Positive slope, crossing the y axis at 9 and the x axis at 3

OpenStudy (anonymous):

thank you

OpenStudy (lasttccasey):

@mathmale if the equation is f(x) = (x + 3)3 -> f(x) = 3x+9, where did you come up with f(x) = 1/3x + 1? Just curious but that doesn't appear to be a legal mathematical move unless I missed something.

OpenStudy (mathmale):

How embarrassing! You're right; I'm wrong. f(x) = (x+3)3 could also be written as f(x) = 3(x+3). Assuming that this assumption (which is what it is) is correct, then f(x) = 3(x+3) = 3x + 9. What is your slope here? What is your y-intercept?